# Ctrl-Adapter: An Efficient and Versatile Framework for Adapting Diverse Controls to Any Diffusion Model

### v1

@@ D.2 Trade-off between Visual Quality and Spatial Control
In Fig. 21, Fig. 22, and Fig. 23, we show the visual quality (FID) and spatial control (SSIM/optical flow error) metrics with different numbers of denoising steps with spatial control (with the fusion of Ctrl-Adapter outputs) on SDXL, SVD, and I2VGen-XL backbones respectively. Specifically, suppose we use $N$ denoising steps during inference, a control guidance level of $x\%$ means that we fuse Ctrl-Adapter features to the video diffusion UNet during the first $x\%\times N$ denoising steps, followed by $(100-x)\%\times N$ regular denoising steps. In all experiments, we find that increasing the number of denoising steps with spatial control improves the spatial control accuracies (SSIM/optical flow error) but hurts visual quality (FID).
@@
Table 6: Comparison of different weighting methods for multi-condition video generation (see Fig. 24 for visualization of the weighting methods). The control sources are abbreviated as D (depth map), C (canny edge), N (surface normal), S (softedge), Seg (semantic segmentation map), L (line art), and P (human pose).
Conditions Equal Weights Linear Weights MLP Router
(Default) Timestep Text/Image Emb Timestep+Text/Image Emb
FID ($\downarrow$) Flow Error ($\downarrow$) FID ($\downarrow$) Flow Error ($\downarrow$) FID ($\downarrow$) Flow Error ($\downarrow$) FID ($\downarrow$) Flow Error ($\downarrow$) FID ($\downarrow$) Flow Error ($\downarrow$)
D + C 8.50 2.84 9.14 2.89 9.41 3.51 8.73 3.16 8.64 3.31
D + P 11.32 3.48 10.98 3.32 11.13 3.35 11.35 3.37 10.69 3.43
D + C + N + S 8.75 2.40 8.39 2.23 9.51 2.78 7.91 2.76 8.09 2.69
D + C + N + S + Seg + L + P 9.48 2.93 8.18 2.48 8.17 2.45 8.83 2.48 8.51 2.43
Figure 24: Visualization of different routing methods for combining multiple ControlNet outputs.
D.3 Different weighing modules for multi-condition generation
For multi-condition generation described in Sec. 3.3, in addition to simple linear weights, we also experimented with learning a router module that takes additional inputs such as diffusion time steps and image/text embeddings and outputs weights for different ControlNets. We compare (a) equal weights; (b) linear weights, learning a scalar weight to each ControlNet Ctrl-Adapter; (c) MLP router - taking timestep as inputs; (d) MLP router - taking image/text embedding as inputs; and (e) MLP router - taking timestep and image/text embedding as inputs. The MoE router in (c), (d), (e) are constructed as a 3-layer MLP. We illustrate the five methods in Fig. 24.
All weighting schemes perform effectievly.
Table 6 show that all weighting schemes for fusing different ControlNet outputs perform effectively, and no specific method outperforms other methods with significant margins in all settings. Thus, we use a single linear weighting scheme as our default configuration for multi-condition generation.
Figure 25: Linear weights allocated to each control condition in each of the middle and output blocks of Ctrl-Adapter on the I2VGen-XL backbone.
Visualization of learned weight for different conditions.
In Fig. 25, we illustrate the learned linear weights allocated to each control condition in each of the middle and output blocks of Ctrl-Adapter on the I2VGen-XL backbone.
Appendix E Additional Qualitative Examples
We provide additional qualitative examples in the following.
@@ Appendix E Additional Qualitative Examples
In Fig. 26 and Fig. 27, we show image generation results on COCO val2017 split using depth map and canny edge conditions, respectively.
@@
In Fig. 28 and Fig. 29, we show video generation results on depth map and canny edges extracted from videos from Sora [^10] https://openai.com/sora and videos from the internet, using depth map and canny edge conditions, respectively.
In Fig. 30 and Fig. 31, we show image generation results on prompts from Lexica [^11] https://lexica.art/ using depth map and canny edge conditions, respectively.
Figure 26: Image generation results on COCO val2017 split using depth map as control condition.
Figure 27: Image generation results on COCO val2017 split using canny edge as control condition.
@@
Figure 28: Video generation results with I2VGen-XL using depth map as a control condition. See our project page for the full videos.
@@
Figure 29: Video generation results with I2VGen-XL using canny edge as a control condition. See our project page for the full videos.
Figure 30: Image generation results from Lexica prompts with SDXL using depth map as a control condition.
@@
Figure 31: Image generation results from Lexica prompts with SDXL using canny edge as a control condition.

### v2

@@ F.2 Trade-off between Visual Quality and Spatial Control
In Fig. 19, Fig. 20, and Fig. 21, we show the visual quality (FID) and spatial control (SSIM/optical flow error) metrics with different numbers of denoising steps with spatial control (with the fusion of Ctrl-Adapter outputs) on SDXL, SVD, and I2VGen-XL backbones respectively. Specifically, suppose we use $N$ denoising steps during inference, a control guidance level of $x\%$ means that we fuse Ctrl-Adapter features to the video diffusion U-Net during the first $x\%\times N$ denoising steps, followed by $(100-x)\%\times N$ regular denoising steps. In all experiments, we find that increasing the number of denoising steps with spatial control improves the spatial control accuracies (SSIM/optical flow error) but hurts visual quality (FID).
@@
Appendix G Additional Qualitative Analysis
G.1 Visualization of Spatial Feature Maps
As mentioned in Sec. 3 and Sec. E.2, the spatial features encoded in the U-Net of ControlNets and the DiT blocks are structurally different. We visualize this difference in Fig. 22. For the DiT-based model, we use PixArt-$\alpha$ as a representative. We follow the visualization method mentioned in [72]. Specifically, we first extract the spatial features from different DiT blocks and U-Net middle/output blocks at the last denoising step during inference. For each block, we applied PCA to the extracted features and visualized the top three leading components.
As shown in Fig. 22, almost all 28 DiT blocks capture global and semantic information about the object "cactus". This observation is consistent with the findings in [20]. On the other hand, the U-Net blocks in ControlNet demonstrate a coarse-to-fine pattern as the feature map size increases. This indicates that mapping output blocks A/B of ControlNet to DiT blocks is a better option compared to using middle or output blocks C/D of the ControlNet.
Figure 22: Visualization of spatial feature maps in PixArt-$\alpha$ and canny edge ControlNet. We first extract the spatial features from different DiT blocks and U-Net middle/output blocks at the last denoising step during inference. For each block, we applied PCA to the extracted features and visualized the top three leading components. Almost all 28 DiT blocks capture global and semantic information about the object "cactus", while the U-Net blocks in ControlNet demonstrate a coarse-to-fine pattern as the feature map size increases.
G.2 Fast Training Convergence
In addition to the quantitative results shown in Fig. 2, we provide a more straightforward visualization for SDXL depth ControlNet + Ctrl-Adapter training. The training speed test is performed on 4 A100 80GB GPUs, with a batch size of 1 per GPU. As shown in Fig. 23, for relatively easy examples (i.e., bedroom, sandwich, bus), our Ctrl-Adapter training can converge within 4.5 GPU hours (which is equivalent to around 1.125 hours measured in training clock time). For complex examples and those requiring fine details (i.e., surfing man, group of kids), our Ctrl-Adapter can also converge within around 6 to 7.5 GPU hours (which is equivalent to 1.5 to 1.875 hours measured in training clock time), which proves the training efficiency of our Ctrl-Adapter.
Figure 23: Training efficiency of Ctrl-Adapter on SDXL backbone. Total training GPU hours are measured on 4 A100 80GB GPUs, with batch size per GPU equal to 1.
Appendix H Additional Visualization Examples
We provide more qualitative examples in this section.
H.1 Video Generation Visualization Examples
@@
In Fig. 24, we show video generation results on COCO val2017 split using depth map and canny edge as control conditions. We visualize baseline methods as well as Ctrl-Adapters built on top of Hotshot-XL [49], SVD [3], I2VGen-XL [89], and Latte [46].
@@
In Fig. 25 and Fig. 26, we show video generation results with I2VGen-XL using depth map and canny edge extracted from videos from Sora [^9] https://openai.com/sora and the internet.
In Fig. 27, we show video generation results with Latte using soft edge extracted from videos from Sora [^10] https://openai.com/sora and the internet.
Figure 24: Videos generated from different video control methods and Ctrl-Adapter on DAVIS 2017, using depth map (left) and canny edge (right) conditions.
@@
Figure 25: Video generation with I2VGen-XL + Ctrl-Adapter using depth map as a control condition.
@@
Figure 26: Video generation with I2VGen-XL + Ctrl-Adapter using canny edge as control condition.
Figure 27: Video generation with Latte + Ctrl-Adapter using soft edge as control condition.
H.2 Multi-Condition Video Generation Visualization Examples
Fig. 28 shows example videos generated with single and multiple conditions. While all videos correctly capture the high-level dynamics of ‘a woman wearing purple strolling during sunset’, the videos generated with more conditions show more robustness in several minor artifacts. For example, when only depth map is given (Fig. 28 a), the building behind the person is blurred. When depth map and human pose are given (Fig. 28 b), the color of the purse changes from white to purple. When four conditions (depth map, canny edge, human pose, and semantic segmentation) are given, such artifacts are removed (Fig. 28 c).
In Fig. 29, we show multi-condition control examples with I2VGen-XL.
Figure 28: Video generation from single and multiple conditions with Ctrl-Adapter on I2VGen-XL. (a) single condition: depth map; (b) 2 conditions: depth map + human pose; (c) 4 conditions: depth map + human pose + canny edge + semantic segmentation. Adding more conditions can help fix several minor artifacts (e.g., in (a) – building is blurred; in (b) – purse color changes).
Figure 29: Video generated with I2VGen-XL + Ctrl-Adapter from 4 control conditions: depth map + human pose + canny edge + semantic segmentation map.
H.3 Image Generation Visualization Examples
In Fig. 30 and Fig. 31, we show image generation results on COCO val2017 split using depth map and canny edge as control conditions.
@@ H.3 Image Generation Visualization Examples
In Fig. 32, Fig. 33, and Fig. 34, we show image generation results on prompts from Lexica [^11] https://lexica.art/ using depth map, canny edge, and soft edge as control conditions.
Figure 30: Image generation from different SDXL-based image control methods and Ctrl-Adapter on COCO val2017 split using depth map as control condition.
Figure 31: Image generation from different SDXL-based image control methods and Ctrl-Adapter on COCO val2017 split using canny edge as control condition.
Figure 32: Image generation with SDXL + Ctrl-Adapter using depth map as a control condition.
Figure 33: Image generation with SDXL + Ctrl-Adapter using canny edge as a control condition.
Figure 34: Image generation with PixArt-$\alpha$ + Ctrl-Adapter using soft edge as a control condition.
H.4 Visualization Examples for Additional Downstream Tasks
Here, we describe in detail how our Ctrl-Adapter can be seamlessly integrated into a wide variety of downstream tasks including video editing, video style transfer, and text-guided motion control, as mentioned in Sec. 6. Additional visualizations are also shown in this part.
Video editing.
Video editing can be achieved by combining image and video Ctrl-Adapters. The procedure is as follows:
  • Firstly, given a source video, we first extract the control condition(s). We can either extract a single control condition (e.g., depth map), or multiple control conditions (e.g., depth map, canny edge, segmentation, etc.) to enhance performance (as we observe in Sec. H.2 that multi-condition control improves spatial control accuracy) .
  • Next, given a user-provided prompt together with the extracted control condition(s), we can use image Ctrl-Adapter (i.e., SDXL + Ctrl-Adapter) to generate the first frame of the video.
  • Finally, we can use video Ctrl-Adapter (i.e., I2VGen-XL + Ctrl-Adapter), with the generated first frame image, text prompt, and extracted control conditions as inputs for final video generation.
In Fig. 35, we provide additional visualizations of the camel example in our main paper.
Figure 35: Video editing by combining SDXL and I2VGen-XL, where both models are equipped with spatial control via Ctrl-Adapter. First, we extract conditions (e.g., depth map) from the original video. Next, we create the initial frame with SDXL + Ctrl-Adapter. Lastly, we provide the newly generated initial frame and frame-wise conditions to I2VGen-XL + Ctrl-Adapter to obtain the final edited video. This video editing framework can edit both object and background.
Text-Guided Motion Control.
This task can be achieved by combining video Ctrl-Adapter with inpainting ControlNet. We train such Ctrl-Adapter as follows:
  • Firstly, for each training video, we randomly select a random block in the image, with the width and height of the block uniformly sampled from $0.25$ to $0.75$ of the image size.
  • Next, we color the block area of the video frames as black color (these processed frames can be regarded as control condition sequences like depth maps).
  • Finally, we can train Ctrl-Adapter with the frozen inpainting ControlNet similar as other types of Ctrl-Adapters mentioned in our main paper.
In Fig. 36, we provide additional visualizations of text-guided image amination.
Figure 36: Text-guided motion control by combining inpainting ControlNet with I2VGen-XL + Ctrl-Adapter. Specifically, inpainting ControlNet takes the masked frames as well as text prompt as inputs. The output feature maps of inpainting ControlNet are then given to Ctrl-Adapter built on top of I2VGen-XL to generate the final video. Object(s) in the masked area can follow the motion described in the text prompt. The unmasked area can be either static or dynamic.
Video style transfer.
This task can be achieved by combining video Ctrl-Adapter with shuffle ControlNet. We train such Ctrl-Adapter as follows:
  • Firstly, for the first frame of each training video, we apply the content shuffle detector implemented in the controlnet_aux [^12] https://github.com/huggingface/controlnet_aux library, to get a shuffled image.
  • Next, we repeat this shuffled image $N$ times, with $N$ equal to the number of output frames of the backbone video diffusion model. These repeated images can be regarded as control condition sequences like depth maps.
  • Finally, we can train Ctrl-Adapter with the frozen shuffle ControlNet similar as other types of Ctrl-Adapters mentioned in our main paper.
in Fig. 37, we provide additional visualizations of video style transfer.
Figure 37: Video style transfer by combining shuffle ControlNet with Latte + Ctrl-Adapter. Specifically, shuffle ControlNet takes the shuffled image as well as text prompt as inputs. The output feature maps of shuffle ControlNet are then given to Ctrl-Adapter built on top of Latte to generate the final video. The generated videos maintain similar style as the input image before shuffling.
Appendix I Broader Impacts
Ctrl-Adapter is motivated by the fact that training ControlNets for new diffusion models, especially video diffusion models that need to consider temporal consistency, can be a huge burden for many users. As shown in Fig. 2, by adopting pretrained ControlNets, training Ctrl-Adapter can be significantly faster than training other controllable image or video generation methods. For example, with the same type of compute (i.e., A100 80GB GPUs), Ctrl-Adapter trained on SDXL depth ControlNet for 10 GPU hours can outperform SDXL ControlNet trained for 700 GPU hours. This drastically reduces the carbon emissions footprint by over 70 times. Therefore, we believe that our work can be a strong contribution to efficient and controllable image and video generation.
While our framework can benefit numerous applications in controllable generation, similar to other image and video generation frameworks, it can also be used for potentially harmful purposes (e.g., creating false information or misleading videos). Therefore, it should be used with caution in real-world applications.
Appendix J Limitations
Note that since Ctrl-Adapter is a method to equip current open-source image and video diffusion models with better control, its performance, quality, and potential visual artifacts largely depend on the capabilities (e.g., motion styles and video length) of the backbone models used. For example, if a diffusion model cannot handle complex motions, Ctrl-Adapter built on top of this backbone might lead to non-optimal complex motion control.
Appendix K License
We use standard licenses from the community and provide the following links to the licenses for the datasets, codes, and models that we used in this paper. For further information, please refer to the specific link.
PyTorch [1]: BSD-style
HuggingFace Transformers [79]: Apache License 2.0
HuggingFace Diffusers [73]: Apache License 2.0
ControlNet [86]: Apache License 2.0
SDXL [52]: MiT License
PixArt-$\alpha$ [8]: AGPL-3.0 License
I2VGen-XL [89]: MiT License
Stable Video Diffusion (SVD) [3]: MiT License
Hotshot-XL [49]: Apache License 2.0
LAION dataset [67]: MiT License
Panda70M dataset [9]: License
COCO dataset [42]: CC BY 4.0
DAVIS 2017 dataset [53]: CC BY 4.0